40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 15 00:40:19 2019, gautam, Update, SUS, ETMY diagnosis plan ETMY_sensorSpectra.pdf
    Reply  Wed May 15 10:57:57 2019, gautam, Update, SUS, EY chamber opened 
       Reply  Wed May 15 17:46:24 2019, gautam, Update, SUS, ETMY inspection firstContactFiber.JPGETMY_sensorSpectra.pdf
          Reply  Wed May 15 19:36:29 2019, Koji, Update, SUS, ETMY instepction 
             Reply  Thu May 16 13:07:14 2019, gautam, Update, SUS, First contact residue removal etmy_sensors.pdfetmy_BRmode.pdf
                Reply  Thu May 16 23:31:55 2019, gautam, Update, SUS, ETMY suspension characterization 6x
                   Reply  Fri May 17 17:01:08 2019, gautam, Update, SUS, ETMY suspension characterization etmy_oplevs.pdf
          Reply  Mon May 20 11:33:46 2019, gautam, Update, SUS, ITMY inspection 
             Reply  Mon May 20 17:12:57 2019, gautam, Update, SUS, ETMY LL adjustment 
                Reply  Mon May 20 22:06:07 2019, gautam, Update, SUS, ITMY also kicked 
                   Reply  Tue May 21 00:15:21 2019, gautam, Update, SUS, Main objectives of vent achieved (?) etmy_oplevs_20190520.pdfETMY_cplxTF.pdfETMY_diagComp.pdfScreen_Shot_2019-05-21_at_12.37.08_AM.png
                      Reply  Tue May 21 21:33:27 2019, gautam, Update, SUS, ETMY HR face cleaned 
                         Reply  Wed May 22 11:53:50 2019, gautam, Update, SUS, ETMY EQ stops backed out ETMY_eigenmodes.pdf
Message ID: 14625     Entry time: Mon May 20 17:12:57 2019     In reply to: 14623     Reply to this: 14627
Author: gautam 
Type: Update 
Category: SUS 
Subject: ETMY LL adjustment 

Following the observation that the response in the LL shadow sensor was lower than that of the others, I decided to pull it out a little to move the signal level with nominal DC bias voltage applied was closer to half the open-voltage. I also chose to rotate the SIDE OSEM by ~20 degrees CCW in its holder (viewed from the south side of the EY chamber), to match more closely its position from a photo prior to the haphazhard vent of the summer of 2018. For the SIDE OSEM, the theoretical "best" alignment in order to be insensitive to POS motion is the shadow sensor beam being horizontal - but without some shimming of the OSEM in the holder, I can't get the magnet clear of the teflon inside the OSEM.

While I was inside the chamber, I attempted to minimize the Bounce/Roll mode coupling to the LL and SIDE OSEM channels, by rotating the Coil inside the holder while keeping the shadow sensor voltage at half-light. To monitor the coupling "live", I set up DTT with 0.3 Hz bandwidth and 3 exponentially weighted averages. For the LL coil, I went through pi radians of rotation either side of the equilibrium, but saw no significant change in the coupling - I don't understand why.

In any case, this wasn't the most important objective so I pushed ahead with recovering half-light levels for all the shadow sensors and closed up with the light doors. I kicked the optic again at 1712:14 PDT, let's see what the matrix looks like now.


before starting this work, i had to key the unresponsive c1auxey VME crate.

ELOG V3.1.3-